# Is the age pension in Australia sustainable and fair? Evidence from forecasting the old-age dependency ratio using the Hamilton-Perry model

### v1

@@ 5 Determining sustainable pension age
Figure 11 shows the anticipated population of Australia from 2022 to 2051. The solid line indicates the current population, while the dashed line within the shaded area shows the point forecast population for 2022 to 2051. The grey area represents the 95% prediction interval for the total population. Under the scenario where the population increases rapidly, the target pension age to maintain sustainability will need to be increased.
@@
Figure 11: 30-year forecasts of the total population, along with 95% prediction intervals.
@@
Figure 12 represents a projection graph that provides a 30-year forecast of the pension age scheme, accompanied by 95% prediction intervals. The primary focus of Figure 12 is delineating the minimum pension age from approximately 2022 to 2051. Central to the figure are two key trend lines: the mean, represented by a black solid line, and the median, illustrated by a red dashed line. Surrounding these central trend lines are blue dashed lines depicting the 95% prediction intervals, offering a probabilistic range within which future values are expected to fall. The trend lines exhibit an upward trajectory, suggesting that the pension age is anticipated to increase over the forecast period. The prediction intervals denote the uncertainty inherent in such long-term forecasts, reflecting potential variability due to economic, demographic, or policy changes. The gradual rise in pension age encapsulated in this figure underscores the demographic and fiscal challenges anticipated to influence pension policy.
@@
Figure 12: 30-year forecasts of the pension age scheme with 95% prediction intervals.
@@
Figure 12 emphasizes this forecast’s critical nature in ensuring the sustainability of the Australian pension system, suggesting that policy adjustments (e.g., raising the pension age) may be necessary to maintain the scheme’s long-term viability. Based on this visualization, we can infer that the current pension age of 67 will remain sustainable until 2025. However, if the Australian Government wants the scheme to remain sustainable until around 2051, it must increase the pension age to 69. In the worst-case scenario, raising the pension age to 69.5 before 2051 will be necessary to ensure sustainability. It is essential to note that any increase in the target pension age must be done carefully and thoughtfully.
@@ 6 Pension welfare distribution
We first use the sub-national age-specific mortality rates obtained from AHMD (2023) and the functional time series method to forecast life expectancies for Australians over the period between 2022–2051. Figure 13 indicates that all states and territories of Australia are expected to witness steady increases in residents’ life expectancies.
@@
Figure 14: Maximum basic age pension rates for singles and couples (black line) in Australia between 1971 and 2023. The forecasts of age pension rates until 2051 are represented by blue curves, with the corresponding 95% prediction intervals indicated by the shaded area.
Combining forecasts of life expectancies and the maximum basic pension rates can provide insight into the distribution of age pension welfare in Australia. We consider single and couple residents in each state and territory across Australia who are eligible for the age pension starting from the following years: 2022, 2027, 2032, 2037, 2042, 2047, and 2051. For ease of comparison, Table 1 shows the present values in 2023 of the projected future age pension payments discounted using the average real interest rates between 1971 and 2023 obtained from The World Bank (2023).
@@
From Table 1, NT and TAS residents are expected to receive much lower age pension income throughout their retirement than the remaining six states and territories. The below-average age pension welfare distributed to retirees in NT and TAS underscores the substantial income disparity in these areas, as indicated by the lowest Gini coefficients of NT and TAS among the eight states and territories in Australia (Kennedy et al., 2017). Most Australians nowadays partially or fully rely on the age pension as the primary source of income in retirement. In 2023, around 92% of Australians aged 65 and over who received income support were beneficiaries of the age pension (Australian Institute of Health and Welfare, 2023). The inequality in the age-pension welfare distribution over life long may put the extra cost of living pressures on older Australians in regional areas since the negative effects of increasing pension ages are found to disproportionately affect poorer households (Morris, 2022). Hence, we recommend the Federal Government and governments in regional and rural areas consider providing retirees with low‐socioeconomic status extra financial and health support.
Lifetime age pension projections on an annual basis were also completed for years between 2023 and 2032. The results also indicate that NT and TAS residents will likely receive much lower lifetime age pension income than pensioners in NSW and VIC. Due to space constraints, the annual age-pension projection results are available at https://yy135.shinyapps.io/shiny/.
@@ 7 Conclusion
Our paper proposes a statistical approach to model and forecast sub-national populations by age and sex. Population forecasts at a more granular sub-national level are needed to facilitate local-level decisions, for example, health and regional migration policies. These forecasts can also reveal patterns of vulnerable groups and track the effects of policy responses. By aggregating the sub-national population forecasts to the national level, we measure the effects of changes in the pension age on the OADR, which we take as a crude measure of the financial burden of the age pension. We have applied stochastic models to forecast the future population age structure and then used those future age structure projections to determine the optimal pension age scheme that will ensure a consistent and desirable level of OADR.
@@
Through the research conducted in this paper, it has been demonstrated that the decision to raise the pension age to 67 has had a successful impact on halting the upward trend in OADR values. While this may have provided a short-term solution, more is needed to ensure a stable OADR in the long run, especially considering the continuous changes in the population’s age profile. Therefore, it is suggested that the pension age be regularly reviewed through the proposed statistical approach to maintain a sustainable OADR.
@@
Although the dataset timeframe is until 2021, including two years of data after the pandemic, we have yet to consider the impact of COVID-19 in this analysis since it is too early to estimate its effect on future age profiles accurately. We made no assumptions about mortality using the HP method. It is known that COVID-19 has mainly affected older people’s mortality rates, which should decrease the OADR. However, Australia’s limited number of deaths caused by COVID-19 suggests that the pandemic’s effect is unlikely to alter our recommendations significantly.
@@ Acknowledgment
The authors are grateful for Dr Tom Wilson’s comments. This research is partially funded by the DataX research center at Macquarie University.

### v2

@@ 5 Determining sustainable pension age
Figure 12 shows the anticipated population of Australia from 2022 to 2051. The solid line indicates the current population, while the dashed line within the shaded area shows the point forecast population for 2022 to 2051. The grey area represents the 95% prediction interval for the total population. Under the scenario where the population increases rapidly, the target pension age to maintain sustainability will need to be increased. Figure 12 (b) presents the forecast using the MFTS method. Similar to Figure 12 (a), the blue line represents actual population data up to 2021, while the red line shows the forecasted population from 2022 to 2051. The grey shaded area indicates the 95% prediction intervals. Notably, the point forecasts using the MFTS method are slightly lower compared to the univariate method, and the prediction intervals are wider. This indicates that the MFTS method accounts for additional uncertainty by capturing the correlations between male and female populations, which may lead to a more cautious projection.
(a) Total population via the univariate functional time series method
(b) Total population via the multivariate functional time series method
@@
Figure 12: 30-year forecasts of the total population, along with 95% pointwise prediction intervals.
The overall trend in both figures suggests that the population of Australia will continue to grow, although the rate of growth and the degree of uncertainty vary between the two methods. The MFTS method provides a more comprehensive forecast by incorporating the interdependencies between different demographic groups, potentially leading to more robust planning and policy decisions. For instance, under scenarios of rapid population increase, the target pension age necessary to maintain sustainability may need to be adjusted upwards. The MFTS method’s broader prediction intervals highlight the importance of preparing for a range of future outcomes, ensuring that policy measures remain effective under varying demographic conditions.
@@
Figure 13 represents a projection graph that provides a 30-year forecast of the pension age scheme, accompanied by 95% prediction intervals. The primary focus of Figure 13 is delineating the minimum pension age from approximately 2022 to 2051. Central to the figure are two key trend lines: the mean, represented by a black solid line, and the median, illustrated by a red dashed line. Surrounding these central trend lines are blue dashed lines depicting the 95% prediction intervals, offering a probabilistic range within which future values are expected to fall. The trend lines exhibit an upward trajectory, suggesting that the pension age is anticipated to increase over the forecast period. The prediction intervals denote the uncertainty inherent in such long-term forecasts, reflecting potential variability due to economic, demographic, or policy changes. The gradual rise in pension age encapsulated in this figure underscores the demographic and fiscal challenges anticipated to influence pension policy. Figure 13 (b) illustrates the pension age scheme forecast using the MFTS method. Similar to Figure 13 (a), the black solid line represents the mean forecast, the red dashed line indicates the median forecast, and the blue dashed lines show the 95% prediction intervals. Notably, the point forecasts for the MFTS method are slightly higher than those obtained using the univariate method, and the prediction intervals are marginally wider. This indicates that the MFTS method captures additional uncertainty by accounting for the correlations between male and female populations, resulting in a more conservative projection.
(a) Pension age scheme with the female and male population forecasts obtained from the univariate functional time series method
(b) Pension age scheme with the female and male population forecasts obtained from the multivariate functional time series method
@@
Figure 13: The 30-year forecasts of the pension age scheme with 95% pointwise prediction intervals.
@@
Figure 13 emphasizes this forecast’s critical nature in ensuring the sustainability of the Australian pension system, suggesting that policy adjustments (e.g., raising the pension age) may be necessary to maintain the scheme’s long-term viability. Based on this visualization, we can infer that the current pension age of 67 will remain sustainable until 2025. However, if the Australian Government wants the scheme to remain sustainable until around 2051, it must increase the pension age to 69. In the worst-case scenario, raising the pension age to 69.5 before 2051 will be necessary to ensure sustainability. It is essential to note that any increase in the target pension age must be done carefully and thoughtfully. The overall trend in both subfigures of Figure 13 indicates that the minimum pension age will need to rise to maintain the sustainability of the pension system. According to the MFTS method, the optimal pension age in 2051 is projected to be 69.9 years, compared to 69.2 years with the univariate method. Under the worst-case scenario forecasted by the MFTS method, the pension age may need to increase to 70 years to ensure sustainability, compared to 69.5 years with the univariate method.
These forecasts underscore the importance of policy adjustments to address demographic changes and fiscal challenges. Ensuring the long-term viability of the pension system will require careful consideration of these forecasts and the potential need for gradual increases in the pension age. This visualization highlights the critical nature of these adjustments, emphasizing the need for thoughtful and deliberate policy measures to sustain the pension system for future generations.
@@ 6 Pension welfare distribution
We first use the sub-national age-specific mortality rates obtained from AHMD (2023) and the functional time series method to forecast life expectancies for Australians over the period between 2022 and 2051. Figure 14 indicates that all states and territories of Australia are expected to witness steady increases in residents’ life expectancies.
@@
Figure 15: Maximum basic age pension rates for singles and couples (black line) in Australia between 1971 and 2023. The forecasts of age pension rates until 2051 are represented by blue curves, with the corresponding 95% pointwise prediction intervals indicated by the shaded area.
Combining forecasts of life expectancies and the maximum basic pension rates can provide insight into the distribution of age pension welfare in Australia. We consider single and couple residents in each state and territory across Australia who are eligible for the age pension starting from the following years: 2022, 2027, 2032, 2037, 2042, 2047, and 2051. For ease of comparison, Table LABEL:tab:1 shows the present values in 2023 of the projected future age pension payments discounted using the average real interest rates between 1971 and 2023 obtained from The World Bank (2023).
@@
From Table LABEL:tab:1, NT and TAS residents are expected to receive much lower age pension income throughout their retirement than the remaining six states and territories. The below-average age pension welfare distributed to retirees in NT and TAS underscores the substantial income disparity in these areas, as indicated by the lowest Gini coefficients of NT and TAS among the eight states and territories in Australia (Kennedy et al., 2017). Most Australians nowadays partially or fully rely on the age pension as the primary source of income in retirement. In 2023, around 92% of Australians aged 65 and over who received income support were beneficiaries of the age pension (Australian Institute of Health and Welfare, 2023). The inequality in the age-pension welfare distribution over life long may put the extra cost of living pressures on older Australians in regional areas since the negative effects of increasing pension ages are found to disproportionately affect poorer households (Morris, 2022). Hence, we recommend that the federal government and governments in regional and rural areas consider providing retirees with low socioeconomic status with extra financial and health support.
Lifetime age pension projections were completed on an annual basis for the years between 2023 and 2032. The results also indicate that NT and TAS residents will likely receive much lower lifetime age pension income than pensioners in NSW and VIC. Due to space constraints, the annual age-pension projection results are available at https://yy135.shinyapps.io/shiny/.
@@ 7 Conclusion
Our paper proposes a statistical approach to model and forecast sub-national populations by age and sex. Population forecasts at a more granular sub-national level are needed to facilitate local-level decisions, such as health and regional migration policies. These forecasts can also reveal patterns of vulnerable groups and track the effects of policy responses. By aggregating the sub-national population forecasts to the national level, we measure the effects of changes in the pension age on the OADR, which we take as a crude measure of the financial burden of the age pension. We have applied stochastic models to forecast the future population age structure and then used those future age structure projections to determine the optimal pension age scheme that will ensure a consistent and desirable level of OADR.
@@
Through the research conducted in this paper, it has been demonstrated that the decision to raise the pension age to 67 has had a successful impact on halting the upward trend in OADR values. While this may have provided a short-term solution, more is needed to ensure a stable OADR in the long run, especially considering the continuous changes in the population’s age profile. Therefore, it is suggested that the pension age be regularly reviewed using the proposed statistical approach to maintain a sustainable OADR.
@@
Although the dataset timeframe is until 2021, including two years of data after the pandemic, we have yet to consider the impact of COVID-19 in this analysis since it is too early to estimate its effect on future age profiles accurately. We made no assumptions about mortality using the HP model. It is known that COVID-19 has mainly affected older people’s mortality rates, which should decrease the OADR. However, Australia’s limited number of deaths caused by COVID-19 suggests that the pandemic’s effect is unlikely to alter our recommendations significantly.
In our calculation, we acknowledge that many factors should be considered, and they do not factor into living expenses in different states. In addition, we do not consider the type of job, as such data are not publicly available in Australia. Our projection and findings consider a demographic factor based on the average remaining life expectancy. Living longer does not imply a healthy life; as a result, it is interesting to study the forecast cost of aging care to the government (see, e.g., Choi et al., 2024).
@@ Acknowledgment
The authors would like to acknowledge the insightful comments from two reviewers, which led to a much-improved paper, and are grateful for Dr. Tom Wilson’s comments. This research is funded by an internal grant from the Data Horizons Consilience Research Center at Macquarie University.